A 50 kg boy riding on 120 kg motorbike starts moving at a speed of 40 km/h. How long to cover will it take a distance of 800m. what is the the engine avg force applied by of the motorbike?
1
Expert's answer
2021-09-24T09:26:50-0400

With the average velocity we can calculate the time:


"v=\frac{d}{t} \implies t=\frac{d}{v}=\frac{0.800\, km}{40\,km/h}=0.02\,h=72\,s"


We use the kinetic energy of the total mass moving with the average velocity to find the average force:

"Fd=\frac{1}{2}m_Tv^2 \implies F=\cfrac{m_T}{2d}\times v^2"


Then we substitute and find

"F=\cfrac{m_T}{2d}\times v^2=\frac{(50+120)kg}{2(800\,m)}\times (\frac{40\,km}{h}\times\frac{10^3\,m}{1\,km}\times\frac{1\,h}{3600\,s})^2"


"\implies F=13.117\,N"


In conclusion, it will take 72 s for the boy to cover the distance, and the average force applied by the engine will be 13.12 N.



Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.